OpenStudy (anonymous):

Simplify the expression. 24times(-1/4)divided(-3) A. –18 B. –2 C. 2 D. 18

OpenStudy (kendricklamar2014):

B

OpenStudy (haleyelizabeth2017):

It is actually C

OpenStudy (kendricklamar2014):

Nope, its B

OpenStudy (haleyelizabeth2017):

-24/4=-6 -6/-3=2

OpenStudy (haleyelizabeth2017):

You forgot that whenever you divide or multiply 2 negative numbers, you get a positive answer

OpenStudy (kendricklamar2014):

24 * (-1/4) = -6

OpenStudy (kendricklamar2014):

-6 / 3 = -2

OpenStudy (haleyelizabeth2017):

It is divided by -3

OpenStudy (haleyelizabeth2017):

24times(-1/4)divided(\(\large-3\))

OpenStudy (kendricklamar2014):

Oh yeah, sorry

OpenStudy (haleyelizabeth2017):

haha it's okay....I do that sometimes :)

OpenStudy (anonymous):

so C

OpenStudy (kendricklamar2014):

Yes

OpenStudy (anonymous):

k tnx

OpenStudy (haleyelizabeth2017):

You are welcome :)